# Intersection and Union Hierarchies of Deterministic

### v2

@@ 6.1 Grammars and Grammatical Trees
(a) Assume that $\hat{x}\sqsubseteq x^{(-)}$. In this case, we obtain $f,g\sqsubseteq x^{(-)}$. Since $c_{i}\prec^{+}c_{s_{i_{0}}}$, we apply Lemma 6.1 and then obtain $s_{i}=t_{i}$ and $^{(s_{i})}T_{w}\equiv{}^{(t_{i})}T_{w^{\prime}}$. From this fact, hereafter, we can focus only on $^{(s_{i})}T_{w}$. Assume that $(v_{1},v_{2})$ generates $(\beta_{1},\beta_{2})$ in $T_{w}$ for two certain sequences $\beta_{1}$ and $\beta_{2}$ of leaves in $T_{w}$. Moreover, assume that $(r_{1},v_{1})$ generates $(\alpha_{1},\alpha_{2})$ in $T_{w}$ for two certain sequences $\alpha_{1}$ and $\alpha_{2}$ of leaves. It then follows from $T_{w}=T_{w}(r_{1})$ that $leaf(T_{w})=\alpha_{1}\cdot leaf(T_{w}(v_{1}))\cdot\alpha_{2}$ and $leaf(T_{w}(v_{1}))=\beta_{1}\cdot leaf(T_{w}(v_{2}))\cdot\beta_{2}$. Notice that $\rho_{1}(\beta_{1})=f$ and $\rho_{1}(\beta_{2})=g$. We then set $x_{1}=\rho_{1}(\alpha_{1})$, $x_{2}=f$, $x_{3}=\rho_{1}(leaf(T_{w}(v_{2}))$, $x_{4}=g$, and $x_{5}y=\rho_{1}(\alpha_{2})$. It is easy to check that the condition (1) is fulfilled.
@@
We then focus on the last (k+1)(h1)+1(k+1)(h-1)+1 subtrees in $^{(s_{i})}T_{w}$. Consider a path $p_{i}$ from $r_{1}$ to $c_{s_{i}+1}$. Each internal node on this path $p_{i}$ has at most $h$ children and it is a parent of one terminal node since $G$ is in Greibach normal form. We claim that the path $p_{i}$ has at least $k+1$ nodes. This is because, otherwise, each internal node on the path $p_{i}$ has at most h1h-1 direct descendants, and thus $^{(s_{i})}T_{w}$ has at most (k+1)(h1)(k+1)(h-1) subtrees, leading to a contradiction. Since $k=|N|$, the pigeonhole principle concludes that there are two nodes, say, $v_{1}$ and $v_{2}$ on the path $p_{i}$ having the same labels. Let $st[v_{1},v_{2}]=(f,g)$. Any node lying between $v_{1}$ and $v_{2}$ on the path $p_{i}$ is a root of a subtree of at most $2^{k\log{h}}$ leaves. From this fact, we conclude that $|f|$ is upper-bounded by $(k+1)2^{k\log{h}}$, and therefore $|f|\leq c$ follows. Note that $f\sqsubseteq\hat{x}$ and $g\sqsubseteq\hat{y}$.
@@
The other cases of $st[v^{\prime}_{1},v^{\prime}_{2}]=(x_{6},u^{\prime}_{2})$ and $(x_{7},u^{\prime}_{2})$ are similarly handled.

### v3

@@ 6.1 Grammars and Grammatical Trees
(a) Assume that $\hat{x}\sqsubseteq x^{(-)}$. In this case, we obtain $f,g\sqsubseteq x^{(-)}$. Since $c_{i}\prec^{+}c_{s_{i_{0}}}$, we apply Lemma 6.1 and then obtain $s_{i}=t_{i}$ and $^{(s_{i})}T_{w}\equiv{}^{(t_{i})}T_{w^{\prime}}$. From this fact, hereafter, we can focus only on $^{(s_{i})}T_{w}$. Assume that $(v_{1},v_{2})$ generates $(\beta_{1},\beta_{2})$ in $T_{w}$ for two certain sequences $\beta_{1}$ and $\beta_{2}$ of leaves in $T_{w}$. Moreover, assume that $(r_{1},v_{1})$ generates $(\alpha_{1},\alpha_{2})$ in $T_{w}$ for two certain sequences $\alpha_{1}$ and $\alpha_{2}$ of leaves. It then follows from $T_{w}=T_{w}(r_{1})$ that $leaf(T_{w})=\alpha_{1}\cdot leaf(T_{w}(v_{1}))\cdot\alpha_{2}$ and $leaf(T_{w}(v_{1}))=\beta_{1}\cdot leaf(T_{w}(v_{2}))\cdot\beta_{2}$. Notice that $\rho_{1}(\beta_{1})=f$ and $\rho_{1}(\beta_{2})=g$. We then set $x_{1}=\rho_{1}(\alpha_{1})$, $x_{2}=f$, $x_{3}=\rho_{1}(leaf(T_{w}(v_{2}))$, $x_{4}=g$, and $x_{5}y=\rho_{1}(\alpha_{2})$. It is easy to check that the condition (1) of the Lemma 3.3 is fulfilled.
@@
We then focus on the last (k+1)(h1)+1(k+1)(h-1)+1 subtrees in $^{(s_{i})}T_{w}$. Consider the path $p_{i}$ from $r_{1}$ to $c_{s_{i}+1}$. Each internal node on this path $p_{i}$ has at most $h$ children and it is a parent of one terminal node since $G$ is in Greibach normal form. We claim that the path $p_{i}$ has at least $k+1$ nodes. This is because, otherwise, each internal node on the path $p_{i}$ has at most h1h-1 direct descendants, and thus $^{(s_{i})}T_{w}$ has at most (k+1)(h1)(k+1)(h-1) subtrees, leading to a contradiction. Since $k=|N|$, the pigeonhole principle concludes that there are two nodes, say, $v_{1}$ and $v_{2}$ on the path $p_{i}$ having the same labels. Let $st[v_{1},v_{2}]=(f,g)$. Any node lying between $v_{1}$ and $v_{2}$ on the path $p_{i}$ is a root of a subtree of at most $2^{k\log{h}}$ leaves. From this fact, we conclude that $|f|$ is upper-bounded by $(k+1)2^{k\log{h}}$, and therefore $|f|\leq c$ follows. Note that $f\sqsubseteq\hat{x}$ and $g\sqsubseteq\hat{y}$.
@@
The other cases of $st[v^{\prime}_{1},v^{\prime}_{2}]\in\{(x_{6},u^{\prime}_{2}),(x_{7},u^{\prime}_{2})\}$ are similarly handled.